# Individualized and Interpretable Sleep Forecasting via a two-stage Adaptive Spatial-Temporal Model

### v1

@@ Abstract
Sleep quality significantly impacts well-being. Therefore, healthcare providers and individuals need accessible and reliable forecasting tools for preventive interventions. This paper introduces an interpretable, individualized two-stage adaptive spatial-temporal model for predicting sleep quality scores. Our proposed framework combines multi-scale convolutional layers to model spatial interactions across multiple input variables, recurrent layers and attention mechanisms to capture long-term temporal dependencies, and a two-stage domain adaptation strategy to enhance generalization. The first adaptation stage is applied during training to mitigate overfitting on the training set. In the second stage, a source-free test-time adaptation mechanism is employed to adapt the model to new users without requiring labels. We conducted various experiments with five input window sizes (3, 5, 7, 9, and 11 days) and five prediction window sizes (1, 3, 5, 7, and 9 days). Our model consistently outperformed time series forecasting baseline approaches, including Long Short-Term Memory (LSTM), Informer, PatchTST, and TimesNet. The best performance was achieved with a three-day input window and a one-day prediction window, yielding a root mean square error (RMSE) of 0.216. Furthermore, the model demonstrated good predictive performance even for longer forecasting horizons (e.g, with a 0.257 RMSE for a three-day prediction window), highlighting its practical utility for real-world applications. We also conducted an explainability analysis to examine how different features influence sleep quality. These findings proved that the proposed framework offers a robust, adaptive, and explainable solution for personalized sleep forecasting using sparse data from commercial wearable devices.
@@ I Introduction
Sleep plays a critical role in physical and mental health, and poor sleep quality is associated with a range of health risks, including cardiovascular disease, cognitive decline, and depression [1]. Most existing research on sleep assessment has focused on classification of sleep stages or retrospective sleep quality estimation, often relying on controlled laboratory data such as polysomnography (PSG) [2]. Studies with PSG provide high-fidelity insights, but they are costly, time-consuming, and impractical to scale for large populations. Furthermore, most studies do not address the predictive task of forecasting future sleep quality, which is essential for proactive health interventions. Moreover, in data sets that do only rely on a single night of recording, the first-night effect might alter symptoms.
@@
Wearable devices represent a promising alternative to traditional polysomnography for sleep monitoring because they have expanded sleep forecasting applications beyond clinical environments, enabling personalized health monitoring. Wearable sensors such as accelerometers, photoplethysmography, and heart rate monitors embedded in commercial smartwatches like Garmin and Fitbit provide promising alternatives for unobtrusive long-term sleep monitoring, demonstrating significant reliability that makes them valuable tools for both personal health management and scientific research [3, 4]. Using data acquired by acti-watches, Convolutional Neural Networks (CNNs) and Long Short-Term Memory (LSTM) networks, a time-batched version of LSTM-RNN (TB-LSTM), these systems can analyze large volumes of data to achieve an operating characteristic (ROC) curve (AUC) of 0.9449 [5]. Moreover, wearable systems using objective sleep data have outperformed subjective self-reports in forecasting cognitive performance [6]. These innovations facilitate individualized feedback, early identification of sleep disorders [7], and population-level studies crucial for informing public health strategies. For instance, global smartphone-derived data has demonstrated how biological and cultural factors influence sleep patterns [8].
@@
Several validation studies have compared the performance of smartwatches against gold-standard methods like polysomnography (PSG) and actigraphy. Fitbit devices, equipped with advanced sensors including accelerometers, optical heart rate monitors, and SpO2 sensors, have shown strong correlations with PSG in detecting sleep and wake periods, with studies reporting agreement rates over 85% for total sleep time and sleep efficiency [4, 9]. Commercial wearables like Fitbit and Garmin enable large-scale sleep research through affordability, cloud storage, and ease of use. Their non-invasive nature improves compliance over traditional PSG methods for longitudinal studies [10]. While not clinical-grade, these reliable devices advance sleep science and personalized health management. Deep learning models mentioned above [5] have achieved high predictive accuracy, but their inherent ”black-box” nature poses significant limitations in clinical utility and user trust. Especially in healthcare applications, interpretability is crucial for validating predictions, guiding treatment decisions, and ensuring fairness and reliability. Explainable AI (XAI) methodologies aim to bridge this gap by providing transparent insights into model decisions and elucidating the importance of various features [11].
@@
Despite this abundance of wearable-generated data, there is a significant gap in the literature when it comes to forecasting sleep quality or sleep score trajectories, especially using data from commercial devices. In this work, we propose a novel adaptive, individualized, and explainable framework that bridges this gap by forecasting sleep quality scores using multimodal time-series data collected from a commercial smartwatch (Garmin Vivosmart 5). Our model is designed to generalize across users using a two-stage domain adaptation strategy and to operate effectively in real-world settings where labeled data are sparse and data distributions vary between individuals.
@@ II Related Work
Machine learning techniques have substantially improved the accuracy and utility of sleep forecasting. Deep learning models, such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), and attention-based[14] architectures, excel at capturing temporal dependencies and modeling complex interactions in sleep data [15].
@@
Figure 1: Illustration of domain adaptation for data from various subjects and conditions
@@ H Hydration milliliter N/A N/A
Test‐time adaptation (TTA) is an emerging inference‐time strategy that dynamically refines a pre‐trained model using incoming test data to address distributional shifts common in sleep forecasting. Unlike traditional models trained on fixed datasets—which often generalize poorly across age groups, health conditions, or different wearable devices—TTA enables on‐the‐fly adjustment of model parameters, improving personalization and accuracy without full retraining [19]. This is particularly valuable for handling inter‐ and intra‐individual variations in sleep patterns due to lifestyle changes or stress, as well as device heterogeneity arising from diverse data collection algorithms [20]. Moreover, by adapting in real time to environmental factors such as ambient noise or temperature fluctuations, TTA enhances the robustness and reliability of sleep‐quality predictions in real‐world settings [21].
Recent studies have integrated Explainable artificial intelligence (XAI) into sleep forecasting systems. In clinical settings, understanding contributing factors such as heart rate variability or environmental conditions enables clinicians to validate predictions against medical knowledge and make informed treatment decisions. For patients, interpretability increases engagement with behavioral interventions by clearly linking actions (e.g., screen time) to sleep outcomes [22, 23]. Additionally, XAI facilitates model improvement by identifying influential features and addressing biases, ensuring equitable and effective predictions across diverse populations [24]. Thus, XAI transforms sleep forecasting models from black-box predictors into transparent, actionable tools for personalized sleep management. Ahadian et al. [25] developed a trustworthy forecasting system employing Temporal Convolutional Networks (TCN), LSTM, and Temporal Fusion Transformers (TFT), complemented by SHAP-based counterfactual explanations and attention mechanisms. Their research shows that interpretability significantly improves model reliability, particularly for detecting sleep disorders across heterogeneous populations. Similarly, Göktepe-Kavis et al. [26] emphasize the necessity of explainable methods within AI-driven sleep scoring and diagnostic frameworks. They highlight the critical role interpretability plays in mitigating algorithmic biases and fostering clinical trust. Furthermore, Kumar et al. [27] stress the importance of interpretability for achieving standardized, generalizable AI models for sleep disorder detection. They argue that explainability is not only vital for user comprehension but also essential for regulatory compliance and ethical deployment in healthcare. Collectively, these studies highlight a growing consensus that explainability is indispensable in AI-driven sleep forecasting, particularly when transitioning predictive models from controlled laboratory settings to real-world clinical and personal healthcare applications.
@@ III-A Dataset
Participants wore Garmin Vivosmart 5 devices for 10-15 weeks to record accelerations, heart rate, ambient light, and blood oxygen saturation, with instructions to remove the device daily as recommended by the manufacturer. They were recruited from two field labs in northern Netherlands: a weekly walking group supervised by healthcare professionals, and a 10-week personal lifestyle intervention program featuring personalized nutrition, cooking workshops, sports sessions, and mindfulness lessons. Research team members attended weekly walking sessions and visited the intervention site biweekly to provide device assistance, while participants otherwise maintained their regular routines. Of the 17 initial participants, one withdrew after three weeks, leaving 16 female participants with a median age of 55.5 years (IQR: 16.25) and median BMI of 28.22 kg/m² (IQR: 10.74), seven of whom had prior smartwatch experience. All data are maintained in the secure, offline Virtual Research Workspace (VRW) administered by the University of Groningen. Prior to downloading and local analysis, data are preprocessed within the VRW in an offline environment to remove sensitive information. More details of the dataset can be found in our previous work [28].
Figure 2: Boxplot for all features and sleep scores
III-B Anomaly and missing data
@@
III-C Features
This study employs a single daily data point (i.e., sparse data) rather than original high-frequency measurements (one data point every 15 minutes) from Garmin, significantly reducing storage and computational demands. A total of 24 features, which were directly exported from the Garmin Connect App, were used in this study, including total kilocalories, total steps, total distance, highly active seconds, active seconds, moderate-intensity minutes, resting heart rate, minimum average heart rate, maximum average heart rate, average waking respiration value, highest respiration value, lowest respiration value, stress average, deep sleep seconds, light sleep seconds, REM sleep seconds, awake sleep seconds, awake count, average sleep stress, restless moment count, lowest respiration, highest respiration, average respiration, and sleep scores as shown in Table I and Figure 2. The predicting target, sleep score (0-100), is calculated using Firstbeat Analytic [29]. We removed the “Hydration” feature for all participants due to an excessive proportion of missing data because of manual input requirements, thereby ensuring a cleaner and more reliable data set. In addition, the first and last days of each participant’s recordings have been excluded due to insufficient device usage.
Selection Methods
We implement a comprehensive feature selection framework to identify the most relevant physiological and activity features for sleep quality prediction. Starting with 23 potential features encompassing activity metrics, heart rate, and respiration with cardio-respiratory metrics, respiration measurements, and sleep stage durations. We recognize that not all features contribute equally to prediction accuracy. Our framework allows dynamic feature selection through multiple methods, with the capability to reduce the feature set from 23 to 15 most informative features. We design this selection process to occur after data preprocessing but before model training, ensuring that only meaningful features are used for learning sleep quality patterns.
First, we computed Pearson correlation coefficients between each feature and sleep quality scores across all time points. We rank features by absolute correlation values and select those exceeding a threshold of 0.05, ensuring we capture both positive and negative relationships. If fewer than 15 features meet this threshold, we include additional high-correlation features to maintain our target dimensionality. Second, our Recursive Feature Elimination (RFE) approach uses a Random Forest regressor as the base estimator to iteratively remove the least important features. We configure RFE to eliminate one feature at a time until reaching our target of 15 features, providing rankings that reflect non-linear feature importance. By reducing from 24 to 15 features, we achieve approximately 32 percent dimensionality reduction, leading to faster training times, reduced overfitting risk, and improved model interpretability without sacrificing predictive performance.
@@ III-F Evaluation metrics
We use Mean Square Error (MSE), Mean Absolute Error (MAE), and Root Mean Square Error (RMSE) as metrics in this work because they are frequently adopted in forecasting tasks, particularly in deep learning contexts, due to their interpretability and computational simplicity. They are defined as:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|.$
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
@@ III-G Leave-one-out cross-validation
In many machine learning problems involving human participants, data collected from different individuals (participants) may vary significantly due to physiological, behavioral, or demographic factors. One standard approach for evaluating the robustness of a model across different individuals is leave-one-out cross-validation (LOOCV). This approach ensures that for each participant, the model is evaluated on data that remain completely unseen during training, providing a realistic estimate of cross-participant generalization performance. There are 16 distinct participants in this work, each associated with a dataset:
@@
In this work, we split the data by participant. Concretely, for each participant $s\in\{1,2,\ldots,16\}$: Practitioners split the training data (from the $N$ subjects) into a separate training set (N2N-2 participants), a validation set (one participant) and a test set (one participant). For instance,
@@
After performing this procedure for each participant $i$, we obtain a set of metrics on the $i$-th fold: MAE $\{MAE_{1},\ldots,MAE_{N}\}$, MSE $\{MSE_{1},\ldots,MSE_{N}\}$, and RMSE $\{RMSE_{1},\ldots,RMSE_{N}\}$. Leave-one-out cross-validation (LOOCV) is a rigorous strategy for evaluating participant-based machine learning tasks. By training on N2N-2 participants , validating on another participant and testing on the remaining participant, we obtain an unbiased estimate of cross-participant generalization.
Algorithm 1 Our model with Two-Stage Domain Adaptation

### v2

@@ Abstract
Sleep quality impacts well-being. Therefore, healthcare providers and individuals need accessible and reliable forecasting tools for preventive interventions. This paper introduces an interpretable, individualized adaptive spatial-temporal model for predicting sleep quality. We designed a hierarchical architecture, consisting of parallel 1D convolutions with varying kernel sizes and dilated convolution, which extracts multi-resolution temporal patterns—short kernels capture rapid physiological changes, while larger kernels and dilation model slower trends. The extracted features are then refined through channel attention, which learns to emphasize the most predictive variables for each individual, followed by bidirectional Long Short-Term Memory (LSTM) and self-attention that jointly model both local sequential dynamics and global temporal dependencies. Finally, a two-stage adaptation strategy ensures the learned representations transfer effectively to new users. We conducted various experiments with five input window sizes (3, 5, 7, 9, and 11 days) and five prediction window sizes (1, 3, 5, 7, and 9 days). Our model consistently outperformed time series forecasting baseline approaches, including LSTM, Informer, PatchTST, and TimesNet. The best performance was achieved with a three-day input window and a one-day prediction window, yielding a Root Mean Squared Error (RMSE) of 0.216. Furthermore, the model demonstrated good predictive performance even for longer forecasting horizons (e.g, with a 0.257 RMSE for a three-day prediction window), highlighting its practical utility for real-world applications. We also conducted an explainability analysis to examine how different features influence sleep quality. These findings proved that the proposed framework offers a robust, adaptive, and explainable solution for personalized sleep forecasting using sparse data from commercial wearable devices.
@@ I Introduction
Sleep plays a critical role in physical and mental health, and poor sleep quality is associated with a range of health risks, including cardiovascular disease, cognitive decline, and depression [5]. Most existing research on sleep assessment has focused on classification of sleep stages or retrospective sleep quality estimation, often relying on controlled laboratory data such as polysomnography (PSG) [21]. Studies with PSG provide high-fidelity insights, but PSG are costly, time-consuming, and impractical to scale for large populations. Furthermore, most studies do not address the predictive task of forecasting sleep quality, which is essential for proactive health interventions. Moreover, in data sets that do only rely on a single night of recording, the first-night effect might alter symptoms.
@@
Wearable sensors provide an alternative to traditional polysomnography for sleep monitoring because they allow sleep forecasting applications to expand beyond clinical (laboratory) environments, enabling personalized health monitoring. Wearable sensors such as ECG, accelerometers [32], Radar [30], photoplethysmography, and heart rate monitors embedded in commercial devices like smartwatches, Garmin, and Fitbit provide promising alternatives for unobtrusive long-term sleep monitoring. Commercial smartwatches demonstrate reliability for basic sleep/wake detection with an average accuracy of 87.2% across 53 devices, consistently performing in the 85-96% range across multiple studies and maintaining this accuracy in both laboratory and real-world home environments [2]. Fitbit devices demonstrated 87.81% sensitivity in detecting sleep periods and achieved 92.97% sensitivity with 52.44% specificity when operating in 4-stage mode, indicating reliable sleep detection capabilities [13]. These devices have demonstrated reliability that makes them valuable tools for both personal health management and scientific research [2, 13]. Using data acquired by acti-watches, Convolutional Neural Network (CNN) and LSTM networks, a time-batched version of TB-LSTM, these systems, combining wearables and machine learning, can analyze large volumes of data to achieve an 0.892 accuracy [24]. Moreover, systems using objective sleep data have outperformed subjective self-reports in forecasting cognitive performance [29]. These innovations in wearable sensors to collect large amounts of data in daily life environments, in combination with data analysis, facilitate individualized feedback, early identification of sleep disorders [16], and population-level studies crucial for informing public health strategies. For instance, smartphone-derived data has demonstrated how biological and cultural factors influence sleep patterns [28].
@@
Several validation studies have compared the performance of smartwatches against gold-standard methods like PSG and actigraphy. Fitbit devices, equipped with advanced sensors including accelerometers, optical heart rate monitors, and SpO2 sensors, can detect sleep and wake periods, with studies reporting agreement rates over 85% between PSG and smartwatch for total sleep time and sleep efficiency [13, 27]. Commercial wearables like Fitbit and Garmin enable large-scale sleep research through affordability, cloud storage, and ease of use. Their non-invasive nature improves compliance over traditional PSG methods for longitudinal studies [19]. Deep learning models mentioned above [24] applied a ”black-box” nature, which poses significant limitations in clinical utility and user trust. Especially in healthcare applications, interpretability is crucial for validating predictions, guiding treatment decisions, and ensuring fairness and reliability. Explainable artificial intelligence (XAI) methodologies aim to bridge this gap by providing transparent insights into model decisions and elucidating the importance of various input features [25].
@@
Despite this abundance of wearable-generated data, forecasting of sleep quality or sleep scores is not often explored in scientific literature, especially using data from commercial devices. In this work, we propose a novel adaptive, individualized, and explainable framework for forecasting sleep quality scores using multimodal time-series data collected from a commercial smartwatch (Garmin Vivosmart 5). Our model employs a two-stage domain adaptation strategy to generalize across different users and function effectively in real-world environments where labeled data is limited and individual data distributions differ.
@@ II Related Work
Machine learning techniques have substantially improved the accuracy and utility of sleep forecasting. Deep learning models, such as CNN, Recurrent neural network (RNN), and attention-based[18] architectures, CNN-transformer [33], excel at capturing temporal dependencies and modeling complex interactions in sleep data [34].
@@ H Hydration milliliter N/A N/A
Test-Time Adaptation (TTA) is an emerging inference‐time strategy that dynamically refines a pre‐trained model using incoming test data to address distributional shifts common in sleep forecasting. This is particularly valuable for handling inter‐ and intra‐individual variations in sleep patterns due to lifestyle changes or stress, as well as device heterogeneity arising from diverse data collection algorithms [26]. Moreover, by adapting in real time to environmental factors such as ambient noise or temperature fluctuations, TTA enhances the robustness and reliability of sleep‐quality predictions in real‐world settings [10].
Recent studies have integrated XAI into sleep forecasting systems. In clinical settings, understanding contributing factors such as heart rate variability conditions enables clinicians to validate predictions against medical knowledge and make informed treatment decisions. For patients, interpretability increases engagement and adherence with behavioral interventions by clearly linking actions (e.g., screen time) to sleep outcomes [20, 23]. Additionally, XAI facilitates model improvement by identifying influential features and addressing biases, ensuring equitable and effective predictions across diverse populations [11]. Thus, XAI transforms sleep forecasting models from black-box predictors into transparent, actionable tools for personalized sleep management. Ahadian et al. [1] developed a trustworthy forecasting system employing Temporal Convolutional Networks (TCN), LSTM, and Temporal Fusion Transformers (TFT), complemented by SHapley Additive exPlanations (SHAP)-based counterfactual explanations and attention mechanisms. Their research shows that interpretability significantly improves model reliability, particularly for detecting sleep disorders across heterogeneous populations. Similarly, Göktepe-Kavis et al. [4] emphasize the necessity of explainable methods within AI-driven sleep scoring and diagnostic frameworks. They highlight the critical role interpretability plays in mitigating algorithmic biases and fostering clinical trust. Furthermore, Kumar et al. [8] stress the importance of interpretability for achieving standardized, generalizable AI models for sleep disorder detection. They argue that explainability is not only vital for user comprehension but also essential for regulatory compliance and ethical deployment in healthcare. Collectively, these studies highlight a growing consensus that explainability is indispensable in AI-driven sleep forecasting, particularly when transitioning predictive models from controlled laboratory settings to real-world clinical and personal healthcare applications.
@@ III-A Dataset
Participants wore Garmin Vivosmart 5 devices for 10-15 weeks to record accelerations, heart rate, and blood oxygen saturation, with instructions to remove the device daily as recommended by the manufacturer. Participants were recruited from two field labs in northern Netherlands: a weekly walking group supervised by healthcare professionals, and a 10-week personal lifestyle intervention program featuring personalized nutrition, cooking workshops, sports sessions, and mindfulness lessons. Research team members attended weekly walking sessions and visited the intervention site biweekly to provide device assistance, while participants otherwise maintained their regular routines. Of the 17 initial participants, one withdrew after three weeks, leaving 16 female participants with a median age of 55.5 years (IQR: 16.25) and a median Body Mass Index (BMI) of 28.22 kg/m² (IQR: 10.74), seven of whom had prior smartwatch experience. All data are maintained in the secure, offline Virtual Research Workspace (VRW) administered by the University of Groningen. Prior to downloading and local analysis, data were preprocessed within the VRW in an offline environment to remove sensitive information.
Figure 1: Distribution of 23 input features and the sleep score target using violin plots with embedded box plots. Features span five categories: activity, heart rate, respiration, sleep stages, and stress. Blue: input features; Orange: prediction target (sleep score).
III-B Features
This study employs a single daily data point (i.e., sparse data) rather than original high-frequency measurements (one data point every 15 minutes) from Garmin, significantly reducing storage and computational demands. We removed the “Hydration” feature for all participants due to an excessive proportion of missing data because of manual input requirements, thereby ensuring a cleaner and more reliable data set. In addition, the first and last days of each participant’s recordings have been excluded due to insufficient device usage. A total of 24 features, which were directly exported from the Garmin Connect App, were used in this study, including total kilocalories, total steps, total distance, highly active seconds, active seconds, moderate-intensity minutes, resting heart rate, minimum average heart rate, maximum average heart rate, average waking respiration value, highest respiration value, lowest respiration value, stress average, deep sleep seconds, light sleep seconds, REM sleep seconds, awake sleep seconds, awake count, average sleep stress, restless moment count, lowest respiration, highest respiration, average respiration, and sleep scores as shown in Table I and Figure 1. The predicting target, sleep score (0-100), is calculated using Firstbeat Analytic [9].
Selection Methods
We implement a comprehensive feature selection framework to identify the most relevant physiological and activity features for sleep quality prediction. Starting with 23 potential features encompassing activity metrics, heart rate, and respiration with cardio-respiratory metrics, respiration measurements, and sleep stage durations. We recognize that not all features contribute equally to prediction accuracy. Our framework allows dynamic feature selection through multiple methods, with the capability to reduce the feature set from 23 to 15 most informative features. By reducing from 24 to 15 features, we achieve approximately 32 percent dimensionality reduction, leading to faster training times, reduced overfitting risk, and improved model interpretability without sacrificing predictive performance. We design this selection process to occur after data preprocessing but before model training, ensuring that only meaningful features are used for learning sleep quality patterns.
First, we computed Pearson correlation coefficients between each feature and sleep quality scores across all time points. We rank features by absolute correlation values and select those exceeding a threshold of 0.05, ensuring we capture both positive and negative relationships. If fewer than 15 features meet this threshold, we include additional correlated features to maintain our target dimensionality. Second, our Recursive Feature Elimination (RFE) approach uses a Random Forest regressor as the base estimator to iteratively remove the least important features. We configure RFE to eliminate one feature at a time until reaching our target of 15 features, providing rankings that reflect non-linear feature importance. By reducing from 24 to 15 features, we achieve approximately 32 percent dimensionality reduction, leading to faster training times, reduced overfitting risk, and improved model interpretability without sacrificing predictive performance.
III-C Anomaly and missing data
@@ III-F Evaluation metrics
We use Mean Squared Error (MSE), Mean Absolute Error (MAE), and Root Mean Squared Error (RMSE) as metrics in this work because they are frequently adopted in forecasting tasks, particularly in deep learning contexts, due to their interpretability and computational simplicity. They are defined as:
[…]
@@ III-G Leave-one-out cross-validation
In many machine learning problems involving human participants, data collected from different individuals may vary significantly due to physiological, behavioral, or demographic factors. One standard approach for evaluating the robustness of a model across different individuals is Leave-one-out Cross-Validation (LOOCV). This approach ensures that for each participant, the model is evaluated on data that remain completely unseen during training, providing a realistic estimate of cross-participant generalization performance. There are 16 distinct participants in this work, each associated with a dataset:
@@
In this work, we split the data by participant. Concretely, for each participant $s\in\{1,2,\ldots,16\}$: Practitioners split the training data (from the $N$ subjects) into a separate training set (N2N-2 participants), a validation set (one participant), and a test set (one participant). For instance,
@@
After performing this procedure for each participant $i$, we obtain a set of metrics on the $i$-th fold: MAE $\{MAE_{1},\ldots,MAE_{N}\}$, MSE $\{MSE_{1},\ldots,MSE_{N}\}$, and RMSE $\{RMSE_{1},\ldots,RMSE_{N}\}$. LOOCV is a rigorous strategy for evaluating participant-based machine learning tasks. By training on N2N-2 participants, validating on another participant, and testing on the remaining participant, we obtain an unbiased estimate of cross-participant generalization.
Algorithm 1 Our Adaptive Multi-scale Temporal Framework